# LMI Optimization Based Multirate Steady-State Kalman Filter Design

### v1

@@ 6.5 Multi-objective Design Results
To demonstrate the multi-objective design capabilities of the LMI framework, we present results for both Kalman filter with pole placement and Kalman filter with $l_{2}$-induced norm constraint.
@@ 6.5.1 Kalman Filter with Pole Placement
Using the pole placement constraint (49), we investigate the trade-off between Kalman filter performance (trace minimization) and convergence rate (eigenvalue constraint $|\lambda|<\bar{r}$). Table 2 and Fig. 4 summarize the results.
[…]
6.5.2 Kalman Filter with $l_{2}$-induced Norm Constraint
Using the $l_{2}$-induced norm constraint (53), we investigate the trade-off between average performance (trace minimization) and worst-case robustness ($l_{2}$-induced norm). $C_{z}=\sqrt{0.1}I_{Nn}$ is selected. The minimum achievable $l_{2}$-induced norm is $\gamma_{\text{opt}}=1.0214$.
[…]
6.5.3 Comparison of Optimal Gains
Table 4 compares the periodic Kalman gains $L_{0}$ (at GPS + wheel speed update) for different design criteria. The $l_{2}$-constrained design ($\bar{\gamma}=1.5\gamma_{\text{opt}}$) yields gains similar to the optimal Kalman filter, while pole placement with $\bar{r}=0.60$ requires larger gains for faster convergence.
@@ 6.6 Design Guidelines
Based on the numerical results, we recommend:
  1. 1.
    Standard applications: Use optimal Kalman filter design for best average performance when model uncertainty is low.
  2. 2.
    Fast response required: Add pole placement constraint with $\bar{r}=0.7$$0.8$ for guaranteed convergence rate with moderate performance penalty.
  3. 3.
    Robustness critical: Add $l_{2}$-induced norm constraint with $\bar{\gamma}=1.5$$2.0\times\gamma_{\text{opt}}$ to limit worst-case error amplification.
  4. 4.
    Combined requirements: Use the multi-objective formulation (55) with appropriate weights to balance all criteria.
The LMI-based approach provides a unified framework for incorporating these diverse requirements within a single convex optimization problem, enabling systematic design of multirate Kalman filters tailored to specific application needs. The MATLAB implementation of the proposed LMI-based multirate Kalman filter design is available at [29].
@@ 7 Conclusion
Numerical validation using state estimation in an automotive navigation system with GPS (1 Hz) and wheel speed (10 Hz) sensors confirmed that the proposed filter achieves position RMSE of 0.600 m and velocity RMSE of 0.268 m/s, well below the raw measurement noise levels. The trade-off analysis demonstrated that the LMI framework provides systematic tuning between convergence rate and estimation accuracy.
The framework is applicable to general linear time-invariant systems with periodic measurement patterns common in robotics, aerospace, and autonomous vehicles. Future directions include extensions to nonlinear systems via extended Kalman filtering, adaptive noise covariance estimation, and distributed sensor networks with asynchronous measurements.
@@ Acknowledgment
The authors acknowledge the use of Claude (Anthropic) for improving the readability and language clarity of the manuscript, and for technical verification of mathematical expressions. Specifically, Claude was used to refine the English expression in the Introduction and Literature Review, and to check dimensional consistency and numerical accuracy of equations. All technical content, research methodology, results, and conclusions are entirely the authors’ own work.

### v3

@@ 6.5 Multi-objective Design Results
To demonstrate the multi-objective design capabilities of the LMI framework, I present results for both Kalman filter with pole placement and Kalman filter with $l_{2}$-induced norm constraint.
@@ 6.5.1 Kalman Filter with Pole Placement
Using the pole placement constraint (54), I investigate the trade-off between Kalman filter performance (trace minimization) and convergence rate (eigenvalue constraint $|\lambda|<\bar{r}$). Table 3 and Fig. 4 summarize the results.
[…]
6.5.2 Kalman Filter with $l_{2}$-induced Norm Constraint
Using the $l_{2}$-induced norm constraint (58), I investigate the trade-off between average performance (trace minimization) and worst-case robustness ($l_{2}$-induced norm). $C_{z}=\sqrt{0.1}I_{n}$ is selected. The minimum achievable $l_{2}$-induced norm is $\gamma_{\text{opt}}=1.0214$.
[…]
6.5.3 Comparison of Optimal Gains
Table 5 compares the periodic Kalman gains $L_{0}$ (at GPS + wheel speed update) for the optimal Kalman filter and the optimal $l_{2}$-induced norm design [20]. The $l_{2}$-optimal design yields a position gain of $1.000$ (compared to $0.283$ for the Kalman filter), indicating that worst-case robustness optimization prioritizes GPS measurements more aggressively, while the velocity and acceleration gains also increase to achieve a smaller maximum eigenvalue magnitude ($0.653$ vs. $0.967$).
@@ 6.6 Design Guidelines
Based on the numerical results, the following design strategy is recommended. For standard applications, the optimal Kalman filter design (Section 4) provides the best average estimation performance under stochastic disturbances characterized by the noise covariances $Q$ and $R$. When fast convergence is required, adding a pole placement constraint with $\bar{r}=0.7$$0.8$ guarantees a minimum decay rate at the cost of a moderate increase in the estimation error covariance. When limiting worst-case performance degradation is critical, the $l_{2}$-induced norm constraint with $\bar{\gamma}=1.5$$2.0\times\gamma_{\text{opt}}$ bounds the worst-case gain from disturbances $(\check{d}_{w},\check{d}_{v})$ to the estimation error. When multiple requirements must be satisfied simultaneously, the combined formulation (60) provides a systematic means of balancing all criteria within a single convex optimization. For robustness against uncertain system matrices, the polytopic uncertainty extension discussed in Remark 5.5 is the appropriate approach.
The LMI-based approach provides a unified framework for incorporating these diverse requirements within a single convex optimization problem, enabling systematic design of multirate Kalman filters tailored to specific application needs. The MATLAB and Python implementation of the proposed LMI-based multirate Kalman filter design is available at [28].
@@ 7 Conclusion
Numerical validation using state estimation in an automotive navigation system with GPS (1 Hz) and wheel speed (10 Hz) sensors, based on Monte Carlo simulation with 500 independent noise realizations, confirmed that the proposed filter achieves a mean position RMSE of 0.561 m, well below the GPS noise standard deviation of 1.0 m (ratio 0.56). The theoretical steady-state RMSE values, obtained by period-averaging the diagonal elements of the $Nn$-dimensional error covariance upper bound $\check{X}^{-1}$ from the LMI solution, exceed the Monte Carlo averages for all states (ratio 1.002–1.046), confirming both the validity and the tightness of the LMI upper bound. The trade-off analysis demonstrated that the LMI framework provides systematic tuning between convergence rate and estimation accuracy.
The framework is applicable in principle to general linear time-invariant systems with deterministic, periodic measurement patterns, where the sensor availability schedule is known a priori and repeats every frame period. The numerical example in this paper serves as a proof-of-concept demonstration with a 3-state system; validation with higher-dimensional systems is an important direction for future work. Other promising directions include robust filter design under polytopic uncertainty (see Remark 5.5 in Section 5), generalization to periodically repeating non-uniform sampling intervals, extensions to descriptor systems [30] and systems with finite delay, sensor configuration switching strategy design guided by the LMI performance metrics, and extensions to nonlinear systems via extended Kalman filtering.
@@ Acknowledgment
The author acknowledges the use of Claude (Anthropic) for improving the readability and language clarity of the manuscript, and for technical verification of mathematical expressions. Specifically, Claude was used to refine the English expression in the Introduction and Literature Review, and to check dimensional consistency and numerical accuracy of equations. All technical content, research methodology, results, and conclusions are entirely the author’s own work.